OpenStudy (anonymous):

If f(x) varies directly with x2, and f(x) = 75 when x = 5, find the value of f(4). A. 100 B. 50 C. 48 D. 25

OpenStudy (anonymous):

Remember that y is equal to f(x). f(x) = k*x^2 75 = k*(5)^2 75 = k*25 75/25 = k 3 = k k = 3

OpenStudy (anonymous):

Now, the equation f(x) = k*x^2 is turned into f(x) = 3x^2.

OpenStudy (decentnabeel):

f(x)=k×x2 f(5)=75=k×52 75=25k find k=3 and so f(x)=3×x2 then f(4)=3×42=3×16

OpenStudy (anonymous):

Plug in x=4 to the equation f(x) = 3x^2, and you'll get the answer.

OpenStudy (anonymous):

ok hold on one sec

OpenStudy (decentnabeel):

answer is option C

OpenStudy (anonymous):

So it's 48. :)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

No problem!